t-3/2t-4 - 1/12=t+1/3t+6
Answer by stanbon(48568)

(t-3)/(2t-4) - 1/12 = (t+1)/(3t+6)
Multiply thru by 12(t^2-4) to get:
8(t-3)(t+2) - (t^2-4) = 4(t+1)(t-2)
8(t^2-t-6) - (t^2-4) = 4(t^2-t-2)
3t^2-4t-36 = 0
t = [4 +- sqrt(16-4*3*-36)]/6
t = [4 +- sqrt448]/6
t = 4.1943 or t = -2.861
==========
cheers,
Stan H.